OpenStudy (anonymous):

simplify: x^2+9x+8/x^2+10x+9

jimthompson5910 (jim_thompson5910):

what do you get when you factor x^2+9x+8

OpenStudy (anonymous):

(x+1) (x+8)

jimthompson5910 (jim_thompson5910):

and when you factor x^2+10x+9 you get ______

OpenStudy (anonymous):

(x+1) (x+9)

jimthompson5910 (jim_thompson5910):

So \[\Large \frac{x^2+9x+8}{x^2+10x+9}\] turns into \[\Large \frac{(x+1)(x+8)}{(x+1)(x+9)}\] The "x+1" terms cancel out (since they divide to 1) leaving you with \[\Large \frac{x+8}{x+9}\] which is your fully simplified final answer.

OpenStudy (anonymous):

yay thanks its right!!

jimthompson5910 (jim_thompson5910):

you're welcome